OpenStudy (austinl):

Solve the following initial value problem. \(y''-2y'+2y=0,~y(\pi)=e^{\pi},~y'(\pi)=0\)

OpenStudy (austinl):

\(r^2-2r+2=0\) \(r=\dfrac{2\pm2i}{2}\) How would you set this up?

OpenStudy (loser66):

r = 1\(\pm\)i

OpenStudy (austinl):

Sorry, I finally worked my way to the answer. \(y(x)=e^x(-\cos(x)+\sin(x))\)

OpenStudy (loser66):

not - cos, just cos

OpenStudy (austinl):

Are you sure about that?

OpenStudy (loser66):

near the bottom of the first page and then continue in the second one

OpenStudy (austinl):

Is that your textbook?

OpenStudy (loser66):

hhahaha... good. I took it last summer,